3(x - 4) -4 = 5x + 6.9 - 3x
Answer by funmath(2933)

3(x - 4) -4 = 5x + 6.9 - 3x
3(x)+3(-4)-4=5x+6.9-3x
3x-12-4=5x-3x+6.9
3x-16=2x+6.9
-2x+3x-16=-2x+2x+6.9
x-16=6.9
x-16+16=6.9+16
x=22.9
Happy Calculating!!!